# ReZero: Boosting MCTS-based Algorithms by Backward-view and Entire-buffer Reanalyze

### v4

@@ 4.2 Theoretical Analysis
For a non-stationary bandit that conforms to the assumptions of Equation 2, denote the total number of rounds as $n$, and the times a sub-optimal arm $i$ is selected in $n$ rounds as $T_{i}(n)$, then selecting actions according to Equation 5 can ensure that $\frac{{\mathbb{E}[T_{i}(n)]}}{n}\to 0$ as $n\to\infty$. Specifically, if we know the $n$ times sample mean $\hat{\mu}^{*}_{n}$ of the optimal arm in advance, then $\mathbb{E}[T_{i}(n)]$ for all sub-optimal arm $i$ satisfies
@@
Otherwise, if we have the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, then for arm $l$,
$\mathbb{E}[T_{l}(n)]\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (7)
@@
$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (8)
where $\Delta_{i}$ is the optimal gap for arm $i$, $\varepsilon$ is a constant in $(0,\Delta_{i})$, and $C$ is the constant in Equation 2.
We provide the complete proof and draw similar conclusions for AlphaZero in Appendix A. Additionally, our method has a lower upper bound for $\mathbb{E}[T_{i}(n)]$. This implies that our algorithm may yield visit distribution more concentrated on the optimal arm. This is a potential worth exploring in future work, especially in offline scenarios [13] where reanalyze becomes the sole method for policy improvement.
@@ 4.3 The ReZero Framework
The technique introduced in Section 4.1 will bring a new problem in practice. The experiments in the Appendix C.3 demonstrate that batching MCTS allows for parallel model inference and data processing, thereby accelerating the average search speed. However, as shown in the Figure 2, to conduct the backward-view reanalyze, we need to divide the batch into $\frac{1}{k+1}$ of its original size. This diminishes the benefits of parallelized search and, on the contrary, makes the algorithm slower.
@@
We propose a new pipeline that is more compatible with the method introduced in Section 4.1. In particular, during the collect phase, we have transitioned from using MCTS to select actions to directly sampling actions based on the policy network’s output. This shift can be interpreted as an alternative method to augment exploration in the collect phase, akin to the noise injection at the root node in MCTS-based algorithms. Figure 10 in Appendix C.3 indicates that this modification does not significantly compromise performance during evaluation. For the reanalyze process, we introduce the periodical entire-buffer reanalyze mechanism. As shown in Figure 3, we reanalyze the whole buffer after a fixed number of training iterations. For each training iteration, we don’t need to run MCTS to reanalyze the mini-batch, only need to sample the mini-batch and execute the gradient descent.
Overall, this design offers two significant advantages: 1 The entire buffer reanalyze is akin to the fixed target net mechanism in DQN [28], maintaining a constant policy target for a certain number of training iterations. Reducing the frequency of policy target updates correspondingly decreases the number of MCTS calls. Concurrently, this does not result in a decrease in performance. 2 Since we no longer invoke MCTS during the collect phase, all MCTS calls are concentrated in the reanalyze process. And during the entire-buffer reanalyze, we are no longer constrained by the size of the mini-batch, allowing us to freely adjust the batch size to leverage the advantages of large batches. Figure 11 shows both excessively large and excessively small batch sizes can lead to a decrease in search speed. We choose the batch size of $2000$ according to the experiment in Appendix C.3.
@@
  • When directly using policy for data collection, action selection is no longer bound by tree search. Thus, previous vectorized environments like [15] can be seamlessly integrated. Besides, this design makes MCTS-based algorithms compatible to existing RL exploration methods like [29].
@@ 5 Experiment
Figure 4: Time-efficiency of ReZero-M vs. MuZero on six representative Atari games and two board games (Connect4, Gomoku). The horizontal axis represents Wall-clock Time (hours), while the vertical axis indicates the Episode Return over 5 episodes. ReZero-M achieves higher time-efficiency than the baseline on most Atari games and Gomoku, and achieve similar time-efficiency on some simpler game like Connect4. Mean of 5 runs, shaded areas are 95% confidence intervals.
Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy. Our primary goal is to improve time efficiency without compromising sample efficiency. To empirically validate the efficiency of our introduced two boosting techniques for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in [30]: MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively. For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces. For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark [30]. The full implementation details available in Appendix B. Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings. In the next subsections, we sequentially explore and try to answer the following three questions:
@@
Atari Board Games
avg. wall time (h) to 100k env. steps $\downarrow$ Pong Breakout MsPacman Qbert Seaquest UpNDown Connet4 Gomoku
ReZero-M (ours) $\mathbf{1}\scriptstyle{\pm 0.6}$ $\mathbf{7.7}\scriptstyle{\pm 0.3}$ $\mathbf{3.8}\scriptstyle{\pm 0.2}$ $\mathbf{5.1\scriptstyle{\pm 0.3}}$ $\mathbf{6.1}\scriptstyle{\pm 0.5}$ $\mathbf{6.9}\scriptstyle{\pm 0.5}$ $\mathbf{10}\scriptstyle{\pm 0.8}$ $\mathbf{9\scriptstyle{\pm 0.8}}$
MuZero [10] $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
Table 1: Average wall-time of ReZero-M vs. MuZero on various tasks. (left) Six Atari games, (right) two board games. The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs.
@@ 5.1 Time Efficiency
Setup In this section, we aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level. In order to facilitate a fair comparison in terms of wall-time, for each run, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless otherwise specified cases). Key parameters are aligned with those from original papers. Specifically, we set the replay ratio (the ratio between environment steps and training steps) [31, 32] to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) [13] is set to 1. For detailed hyper-paramater configurations, please refer to the Appendix B.
Results Figure 4 illustrates the performance comparison in terms of wall-clock time between ReZero-M and the original MuZero across a variety of RL environments. It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks. Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 7 and 8. For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1. The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
Figure 5: The effect of Reanalyze Frequency in ReZero-M on the MsPacman game. The proper reanalyze frequency can improve time and sample efficiency while obtaining the comparable return with MuZero (reanalyze ratio=1).
@@ 5.2 Effect of Reanalyze Frequency
In next two sub-sectons, we will delve into details to explain how ReZero works. Firstly, we will analyze how the entire-buffer reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency. Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$. The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline. Figure 5 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
@@ 5.3 Effect of Backward-view Reanalyze
Table 2: Comparisons about detailed time indicators between MuZero and ReZero-M inside the tree search.
To further validate and understand the advantages and significance of the backward-view reanalyze technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2. The number of function calls is the cumulative value of 100 training iterations on Pong. Avg. time is the average time of a MCTS across all calls. Comparative analysis between ReZero-M and MuZero reveals that the backward-view reanalyze technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations. Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains. Additionally, Theorem 1 demonstrates that the backward-view reanalyze can reduce the regret upper bound, which indicates a better search result. Figure 9 in Appendix C.3 shows a comparison of sample efficiency between using backward-view reanalyze and origin reanalyze process in MsPacman. The experimental results reveal that our method not only enhances the speed of individual searches but also improves sample efficiency. This aligns with the theoretical analysis.
@@ 6 Conclusion and Limitation
In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms. Unlike most existing works, we incorporates information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency. Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains. However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks. Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X [33], constitutes a fertile avenue for future research. These explorations could broadening the application horizons of MCTS-based algorithms.

### v5

@@ 4.2 Theoretical Analysis
For a non-stationary bandit that conforms to the assumptions of Equation 2, denote the total number of rounds as $n$, the prior score for arm $i$ as $P_{i}$, and the number of times a sub-optimal arm $i$ is selected in $n$ rounds as $T_{i}(n)$, then use a sampled estimation instead of UCB value to evaluate a specific arm(like we do in Equation 5) can ensure that $\frac{{\mathbb{E}[T_{i}(n)]}}{n}\to 0$ as $n\to\infty$. Specifically, if we know the $n$ times sample mean $\hat{\mu}^{*}$ of the optimal arm in advance, then $\mathbb{E}[T_{i}(n)]$ for all sub-optimal arm $i$ satisfies
@@
Otherwise, if we have the $n$ times sample mean $\hat{\mu}_{l}$ of a sub-optimal arm $l$, then for arm $l$,
$\mathbb{E}[T_{l}(n)]\leq 1+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2C^{6}}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (7)
@@
$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2C^{6}}}{\varepsilon^{4}P_{1}^{2}}$ (8)
where $\Delta_{i}$ is the optimal gap for arm $i$, $\varepsilon$ is a constant in $(0,\Delta_{i})$, and $C$ is the constant in Eq. 2.
We provide the complete proof and draw similar conclusions for AlphaZero in Appendix A. Additionally, our method has a lower upper bound for $\mathbb{E}[T_{i}(n)]$. This implies that our algorithm may yield visit distribution more concentrated on the optimal arm. This is a potential worth exploring in future work, especially in offline scenarios (Schrittwieser et al., 2021) where reanalyze becomes the sole method for policy improvement.
@@ 4.3 The ReZero Framework
The technique introduced in Section 4.1 will bring a new problem in practice. The experiments in the Appendix D.3 demonstrate that batching MCTS allows for parallel model inference and data processing, thereby accelerating the average search speed. However, as shown in the Figure 2, to conduct the backward-view reanalyze, we need to divide the batch into $\frac{1}{k+1}$ of its original size. This diminishes the benefits of parallelized search and, on the contrary, makes the algorithm slower.
@@
We propose a new pipeline that is more compatible with the method introduced in Section 4.1. In particular, during the collect phase, we have transitioned from using MCTS to select actions to directly sampling actions based on the policy network’s output. This shift can be interpreted as an alternative method to augment exploration in the collect phase, akin to the noise injection at the root node in MCTS-based algorithms. Figure 11 in Appendix D.3 indicates that this modification does not significantly compromise performance during evaluation. For the reanalyze process, we introduce the periodical entire-buffer reanalyze. As shown in Figure 3, we reanalyze the whole buffer after a fixed number of training iterations. For each iteration, we do not need to run MCTS to reanalyze the mini-batch, only need to sample the mini-batch and execute the gradient descent.
Overall, this design offers two significant advantages: 1 The entire buffer reanalyze is akin to the fixed target net mechanism in DQN (Mnih et al., 2013), maintaining a constant policy target for a certain number of training iterations. Reducing the frequency of policy target updates correspondingly decreases the number of MCTS calls. Concurrently, this does not result in a decrease in performance. 2 Since we no longer invoke MCTS during the collect phase, all MCTS calls are concentrated in the reanalyze process. And during the entire-buffer reanalyze, we are no longer constrained by the size of the mini-batch, allowing us to freely adjust the batch size to leverage the advantages of large batches. Figure 12 shows both excessively large and excessively small batch sizes can lead to a decrease in search speed. We choose the batch size of $2000$ according to the experiment in Appendix D.3.
@@
  • When directly using policy for data collection, action selection is no longer bound by tree search. Thus, previous vectorized environments like Weng et al. (2022) can be seamlessly integrated. Besides, this design makes MCTS-based algorithms compatible with existing RL exploration methods like Badia et al. (2020).
@@ 5 Experiment
Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy. Our primary goal is to improve time efficiency without compromising sample efficiency. Here we first give a toy example case in Section 5.1 to intuitively demonstrate the acceleration effect of our design, and the corresponding code example is placed in the Appendix D.4 for readers to quickly understand the algorithm design.
And then, to validate the efficiency of our introduced two boosting techniques for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in Niu et al. (2023): MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively. For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline. It is important to highlight that ReZero can be seamlessly integrated with various algorithm variants of MuZero. In this context, we have chosen to exemplify this integration using the above two instances. In order to validate the applicability of our method across various decision-making environments, we opt for 26 representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces, and two continuous control tasks in DMControl (Tunyasuvunakool et al., 2020). For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark (Niu et al., 2023). The full implementation details available in Appendix C. Besides, we emphasize that to ensure a fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings. In Section 5.2, Section 5.3 and Section 5.4, we sequentially explore and try to answer the following three questions:
@@
5.1 Toy Case
Figure 4: Acceleration effect on the toy case. Left is a simple maze environment where the agent starts at point A and receives a reward of size $1$ upon reaching the end point G. Middle shows the search time corresponding to each position when set as the root node. Meanwhile, the root node values obtained during the search are preserved. Right shows the corresponding search time when these root node values are used to assist the search. The comparison shows that the search duration is generally reduced. For specific experimental settings and code, please refer to the Appendix D.4.
Intuitively, it is evident that our proposed method can achieve a speed gain because we eliminate the search of a certain subtree, especially when this subtree corresponds to the optimal action (which often implies that the subtree has a larger number of nodes). We conduct an experiment on a toy example case and include the experimental code in the Appendix D.4. This helps to visually illustrate the speed gain achieved by skipping subtree search and allows readers to quickly understand the algorithm design through simple code. As shown in Figure 4 (Left), we implement a simple $7\times 7$ maze environment where the agent starts at point A and receives a reward of $1$ upon reaching point G. We perform an MCTS with each position in the maze as the root node and recorded the search time in Figure 4 (Middle). It can be seen that regions farther from the end point require more time to search (this is related to our simulation settings, see the appendix for details). For comparision, we also performed searches with each position as the root node, but during the search, we used the root node values obtained from the search in Figure 4 (Middle) to evaluate specific actions. The experimental time was recorded in Figure 4 (Right). It can be seen that after eliminating the search of specific subtrees, the search time was generally reduced. This simple result validates the rationality of our algorithm design. In the next section, we will specifically validate the time efficiency of the ReZero framework in diverse decision-making tasks in Section 5.2.
Figure 5: Time-efficiency of ReZero-M vs. MuZero on four representative Atari games, two continuous control tasks of DMControl (ball_in_cup-catch, walker-stand), and two board games (Connect4, Gomoku). The horizontal axis represents Wall-clock Time (hours), while the vertical axis indicates the Episode Return over 5 evaluation episodes. ReZero-M demonstrates superior time-efficiency compared to the baseline across a diverse set of games, encompassing both image and state observations, discrete and continuous actions, and scenarios involving sparse rewards. These figures compute mean of 5 runs, and shaded areas are 95% confidence intervals.
@@ 5.2 Time Efficiency
Setup: We aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level. In order to facilitate a fair comparison in terms of wall-time, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless specified cases). Key parameters are aligned with those from original papers. After each data collection phase, the model is trained for multiple iterations according to the replay ratio, which is denoted as one training epoch. Specifically, we set the replay ratio (the ratio between environment steps and training steps) (D’Oro et al., 2022) to 0.25, and the reanalyze ratio (the ratio between targets computed from the environment and by reanalysing existing data) (Schrittwieser et al., 2021) is set to 1. For detailed hyper-paramater configurations, please refer to the Appendix C.
Results: Our experiments shown in Figure 5 illustrates the training curves and performance comparisons in terms of wall-clock time between ReZero-M and the MuZero across various decision-making environments. Note that on continuous control tasks we use the sampled version of ReZero-M and MuZero Hubert et al. (2021). The data clearly indicates that ReZero-M achieves a significant improvement in time efficiency, attaining a near-optimal policy in significantly less time for these eight diverse decision-making tasks. To provide an alternative perspective on the time cost of training on the same number of environment steps, we also provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1. For a more comprehensive understanding of our findings, we offer the complete table of the 26 Atari games usually used for sample-efficient RL in Appendix D.1. It reveals that, on most games, ReZero require between 2-4 times less wall-clock time per 100k steps compared to the baselines, while maintaining comparable or even superior performance in terms of episode return. Additional results about another algorithm comparison instance ReZero-E and EfficientZero are detailed in Appendix D.2, specifically in Figure 8 and 9. This setting also supports the time efficiency of our proposed ReZero framework.
Atari DMControl Board Games
avg. wall time (h) to 100k env. steps $\downarrow$ Pong Breakout MsPacman Seaquest ball_in_cup-catch walker-stand Connet4 Gomoku
ReZero-M (ours) $\mathbf{1.0}\scriptstyle{\pm 0.1}$ $\mathbf{3.0}\scriptstyle{\pm 0.8}$ $\mathbf{1.4}\scriptstyle{\pm 0.2}$ $\mathbf{1.9\scriptstyle{\pm 0.4}}$ $\mathbf{2.1}\scriptstyle{\pm 0.2}$ $\mathbf{4.3}\scriptstyle{\pm 0.3}$ $\mathbf{5.5}\scriptstyle{\pm 0.6}$ $\mathbf{4.5\scriptstyle{\pm 0.5}}$
MuZero (Schrittwieser et al., 2019) $4.0\scriptstyle{\pm 0.5}$ $4.9\scriptstyle{\pm 1.8}$ $6.9\scriptstyle{\pm 0.3}$ $10.1\scriptstyle{\pm 0.5}$ $5.6\scriptstyle{\pm 0.4}$ $9.5\scriptstyle{\pm 0.6}$ $9.1\scriptstyle{\pm 0.8}$ $15.3\scriptstyle{\pm 1.5}$
Table 1: Average wall-time of ReZero-M vs. MuZero on various tasks. (left) Four Atari games, (middle) two control tasks, (right) two board games. The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs.
Figure 6: The ablation experiment of Reanalyze Frequency in ReZero-M on the Atari MsPacman game. The proper reanalyze frequency can improve time and sample efficiency while obtaining the comparable return with MuZero (reanalyze ratio=1).
@@ 5.3 Effect of Reanalyze Frequency
In next two sub-sectons, we will delve into details to explain how ReZero works. Firstly, we will analyze how the entire-buffer reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency.
Here, we adjust the periodic reanalyze frequency—which determines how often the buffer is reanalyzed during a training epoch—in ReZero for the MsPacman environment. Specifically, we set reanalyze frequency to $\{0,\frac{1}{3},1,2\}$. The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline. Figure 6 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
@@ 5.4 Effect of Backward-view Reanalyze
Table 2: Comparisons about the detailed time cost indicators between MuZero and ReZero-M inside the tree search.
To further validate and understand the advantages and significance of the backward-view reanalyze technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2. The number of function calls is the cumulative value of 100 training iterations on Pong. Avg. time is the average time of a MCTS across all calls. Comparative analysis between ReZero-M and MuZero reveals that the backward-view reanalyze technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations. Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains. The complexity and implementation of the tree search process directly influences the efficiency gains achieved through the backward-view reanalyze technique. As the tree search becomes more intricate and sophisticated, e.g. Sampled MuZero (Hubert et al., 2021), the time savings realized through this method are correspondingly amplified. Additionally, Theorem 1 demonstrates that the backward-view reanalyze can reduce the regret upper bound, which indicates a better search result. Besides, Figure 10 in Appendix D.3 shows a comparison of sample efficiency between using backward-view reanalyze and origin reanalyze process in MsPacman. The experimental results reveal that our method not only enhances the speed of individual searches but also improves sample efficiency. This aligns with the theoretical analysis.
@@ 6 Conclusion and Limitation
In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms. Unlike most existing works, we incorporate information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency. Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains. However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks. Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X (Padalkar et al., 2023), constitutes a fertile avenue for future research. These explorations could broaden the application horizons of MCTS-based algorithms. Additionally, since in offline training scenarios, reanalyze becomes the only means of policy improvement, this makes the acceleration of the reanalyze phase in ReZero even more critical. Moreover, the potential improvement in search results by ReZero may further improve the training result, rather than merely accelerating the training. Therefore, combining ReZero with MuZero Unplugged (Schrittwieser et al., 2021) is a direction worth exploring for building foundation models for decision-making.